Recycle Calculations

\underline{Problem 1}: A simplified flowsheet for the Union Carbide oxo process is given in Figure 1. The reactor operates at a pressure of 200 psia and a temperature of 373 K. The reaction mechanism is as follows:

P L + CO + H_{2} → IBA + NBA → HV

80% P L is converted and IBA/NBA ratio is 0.1. 1% of IBA is converted.

P L  Propylene
CO  Carbon Monoxide
H_{2}  Hydrogen
IBA Isobutyraldehyde
NBA n-butyraldehyde
HV  Heavy compounds
P  Propane

Propane is an inert component that does not take part in the reaction. The feed is available at 1 atm and 298 K and has the following composition:

CO 0.5 kgmol/s
H_{2} 0.5 kgmol/s
P L 0.47 kgmol/s
P  0.03 kgmol/s

• Determine the overall conversion of propylene to n-butyraldehyde for purge rates of 1% and 0.1%. Assume that all the separation steps (distillation) give perfect splits for the components shown.

• How does the propane flowrate change at the reactor inlet when the purge rate goes from 0.1% to 1%?

\underline{Problem 2}: Consider the ammonia process shown in Figure 2. A feed of 20% N_{2}, 78% H_{2}, and 2% CH_{4} is mixed with two recycle streams and enters a reactor. Here, conversion per pass of N_{2} to NH_{3} is 45% according to the reaction:

N_{2} + 3H_{2} → 2NH_{3}

The ammonia product is recovered by flashing the reactor effluent in two stages and recycling the overhead vapor. If the purge fraction is 5% for the high pressure recycle, what is the methane concentration in the reactor feed?

The split fractions for Flash 1 and Flash 2 are given below

ξ_{H_{2}} ξ_{N_{2}} ξ_{CH_{4}} ξ_{NH_{3}}
Flash 1 0.995 0.99 0.99 0.01
Flash 2 1.000 1.00 1.00 0.02

Figure 1: 2 Figure 2:

3